OpenStudy (anonymous):

Find the taylor series of tan(x) with a = 0;

OpenStudy (king):

is this in english?

OpenStudy (anonymous):

Mathematics.

OpenStudy (amistre64):

isnt it just taylor sin over cos?

OpenStudy (amistre64):

cos=1 -sin=0 -cos=-1 sin=0 cos=1 -sin=0 -cos=-1 sin=0 cos=1 cos starts at 1 and are the evens \[cos(x)=1-\frac{x^2}{2!}+\frac{x^4}{4!}-\frac{x^6}{6!}+\frac{x^8}{8!}...\] and sin(x) is the derivative of cos(x) \[cos(x)'=-(-\frac{2x}{2*1!}+\frac{4x^3}{4*3!}-\frac{6x^5}{6*5!}+\frac{8x^7}{8*7!}...\] \[sin(x)=\frac{x}{1!}-\frac{x^3}{3!}+\frac{x^5}{5!}-\frac{x^7}{7!}...\]

OpenStudy (amistre64):

the other way is to try and determine a pattern for tan(x) from its derivatives

OpenStudy (amistre64):

tan sec^2 2 sec^2 tan 4 sec^2 tan^2 + 2sec^4 its gets messy I believe

OpenStudy (amistre64):

the good news tho might be that tan0=0 so all the tan terms cancels out and your left with C sec^2(0) = 1